Jump to main content.


Drake Chemical

EPA Weekly Situation Reports for November 1998

U. S. ARMY CORPS OF ENGINEERS
DRAKE CHEMICAL SUPERFUND SITE

182 MYRTLE STREET
LOCK HAVEN, PA 17745
PHONE 717-748-4505
FAX 717-748-7549

Wednesday, November 25, 1998

FOR DATES: From 7:00 a.m. Wednesday, November 18, 1998 to 7:00 a.m. Wednesday, November 25, 1998.

The Thermal Destruction Facility (TDF) was on-line for 166.6 hours and processed 5937 tons of contaminated soil during the time frame of this report. The TDF was on-line 99.2 % of the time with an average hourly feed rate of 35.6 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated 193,679.0 tons of contaminated soil, which amounts to approximately 66.2 % of the overall contaminated soil on-site.

As indicated above, the average hourly feed rate for the time frame of this report was reduced to approximately 36 tons per hour when one bottom ash analytical report indicated that the thermally treated soil did not meet the project requirements for backfilling. That feed soil had a high moisture content and the analytical report indicated that the concentration of BNA was slightly above the allowable limit. That material was returned to the exclusion zone for further thermal treatment.

Construction activities during the time frame of this report included TDF operations and maintenance, wastewater management, feed preparation, debris separation, lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, placing odor suppressant foam on the excavation face, cleaning and backfilling of rock and concrete debris, and winterization of temperature sensitive components of the TDF. The contractor continued to work on the installation of the lime silo in the exclusion zone. The contractor continued to perform off-site and perimeter air monitoring (PAM) activities.

Excavation continued along the southern end of the excavation face. A substantial amount of demolition debris and fill was unearthed in the excavation area. No drums were encountered during the time frame of this report. One roll-off (a trailer container) of

hazardous waste was shipped off-site during the time frame of this report. The roll-off contained a composite of drum contents that had been excavated from the Drake property during the incineration phase of the project. The chemical waste was shipped off-site to a licensed hazardous waste facility for thermal treatment. Prior to shipment off-site, the contents of the roll-off were tested for reactivity and chemical compatibility.

Related site activities included:

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling. There is an extended maintenance period tentatively scheduled to run from December 7 to December 14, 1998. During this period, the incinerator will be placed in cold shutdown in order to perform maintenance to the internal components of the TDF. In addition to the preventative maintenance, the contractor will be performing the monthly calibrations and interlock checks. Hazardous waste feed is scheduled to resume on December 15, 1998.

PREPARED BY: Steve Gibson for William Werntges, USACE


Wednesday, November 18, 1998

FOR DATES: From 7:00 a.m. Wednesday, November 11, 1998 to 7:00 a.m. Wednesday, November 18, 1998.

The Thermal Destruction Facility (TDF) was on-line for 166.7 hours and processed 6624.2 tons of contaminated soil during the time frame of this report. The TDF was on-line 99.3 % of the time with an average hourly feed rate of 39.7 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated 188,071.8 tons of contaminated soil, which amounts to approximately 64.3 % of the overall contaminated soil on-site.

Construction activities during the time frame of this report included TDF operations and maintenance, wastewater management, feed preparation, debris separation, lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, the placement odor suppressant foam on the excavation face, rock and concrete debris cleaning and backfilling. The contractor continued to work on the installation of the lime silo in the exclusion zone. The contractor continued to perform off-site and perimeter air monitoring (PAM) activities.

Excavation continued along the southern end of the excavation face. A substantial amount of demolition debris and fill was unearthed in the excavation area. No drums were encountered and no containers of hazardous waste were shipped off-site during the time frame of this report. Approximately 20 tons of activated carbon from the feed preparation building's air handling unit were shipped off-site to a carbon regeneration plant for recycling. The carbon material was shipped as non-RCRA and non-DOT regulated waste.

Related site activities included: * A weekly progress meeting held on Tuesday, November 17, 1998 with representatives from the USEPA, USACE, PADEP, and OHM in attendance. * MRI completed their annual audit of the site and off-site air monitoring equipment. J. Cheney (USACE) and Dr. S. Tucker (OHM) were on-site to witness the audit.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling. Several roll-offs of hazardous and non-hazardous debris are scheduled to be shipped off-site to a hazardous waste landfill.

PREPARED BY: Steve Gibson for David Modricker, USACE


Thursday, November 12, 1998

FOR DATES: From 7:00 a.m. Wednesday, November 4, 1998 to 7:00 a.m. Wednesday, November 11, 1998.

The Thermal Destruction Facility (TDF) was on-line for 139.9 hours and processed 5593.8 tons of contaminated soil during the time frame of this report. The TDF was on-line 83.3 % of the time with an average hourly feed rate of 40.0 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated 183,166.7 tons of contaminated soil, which amounts to approximately 63.8 % of the overall contaminated soil on-site. A site survey conducted last week confirmed the quantity of material yet to be treated. Site surveys will be conducted monthly to confirm daily quantity calculations and verify progress. Initial results show the contractor ahead of schedule.

On Monday, November 9, 1998, the incinerator was placed in hot hold in order to perform monthly interlock checks and perform maintenance to the dual feed screws. The planned down time was extended to facilitate repairs to the dry ash drag. The bottom floor plate of the dry ash drag had been damaged by a piece of metal which was dislodged from the kiln. New carbide cutting teeth were welded onto the existing feed screws in order to reduce wear to the screws. In addition to the preventative maintenance performed, the monthly TDF interlock checks were performed in order to test the automatic shutdown systems of the incinerator. There was a planned Thermal Relief Valve (TRV) opening initiated on Monday, November 9, 1998 at 1404 hours. Prior to the TRV opening, feed to the incinerator had been shut down for over five hours and the necessary regulatory authorities had been notified by OHM.

Construction activities during the time frame of this report included TDF operations and maintenance, wastewater management, feed preparation, debris separation, lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, the placement odor suppressant foam on the excavation face, rock and concrete debris cleaning and backfilling. The contractor continued to perform off-site and perimeter air monitoring (PAM) activities. Repairs were completed to the heat trace line connected to PAM station 3.

Excavation continued along the southern end of the excavation face. No drums were encountered and no roll-offs (containers of hazardous waste) were shipped off-site during the time frame of this report.

Related site activities included:

* A weekly progress meeting held on Tuesday, November 10, 1998 with representatives from the USEPA, USACE, PADEP, and OHM in attendance.

* A site tour was given to students from Penn State University.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling.

PREPARED BY: Steve Gibson for David Modricker, USACE


Wednesday, November 4, 1998

FOR DATES: From 7:00 a.m. Wednesday, October 28, 1998 to 7:00 a.m. Wednesday, November 4, 1998.

The Thermal Destruction Facility (TDF) was on-line for 167.0 hours and processed 6683 tons of contaminated soil during the time frame of this report. The TDF was on-line 99.4 % of the time with an average hourly feed rate of 40.0 tons per hour while on-line. The contractor, OHM, continues to work 24 hours a day, 7 days a week operating and maintaining the TDF.

The contractor has successfully treated 174,937.8 tons of contaminated soil, which amounts to approximately 63.2 % of the overall contaminated soil on-site. A site survey was conducted this week to confirm the quantity of material yet to be treated. The percent complete will be adjusted in the near future based on the results of the most recent survey information.

Construction activities during the time frame of this report included TDF operations and maintenance, wastewater management, feed preparation, debris separation, lime treatment of soils, ash handling, ash backfilling, dust control to prevent fugitive emissions, the placement odor suppressant foam on the excavation face, rock and concrete debris cleaning and backfilling. A lime silo was installed on the ash stabilization pad within the exclusion zone.

The contractor also continued to perform off-site and perimeter air monitoring activities. During this reporting period the primary real time perimeter air monitoring system was shut down in order to make repairs to the system's heat trace lines. While the system was down, site air was monitored with hand held instruments at each of the PAM stations for the required intervals. The primary real time air monitoring system is scheduled to be fully functional later this week.

Excavation continued along the southern end of the excavation face. Much of the excavated material consisted of wood and other demolition debris. Four roll-offs of non-hazardous waste were shipped off-site to a hazardous waste landfill during the time frame of this report. Three of the roll-offs contained used Personal Protective Equipment and miscellaneous site debris, and the fourth roll-off contained drum carcasses and overpacks from previously excavated drums.

Related site activities included:

* A weekly progress meeting held on Tuesday, November 3, 1998 with representatives from the USEPA, USACE, PADEP, and OHM in attendance.

* Site tours were given to students from East Lycoming High School and students from Hartwick College.

Major activities for the next week include: Operation of the TDF, soil feed preparation, excavation, off-site and perimeter air monitoring activities, WWTP operations, ash handling, ash backfilling, rock & inorganic debris handling, rock cleaning, and rock backfilling. There is a planned shut down (hot hold) of the TDF scheduled for November 9, 1998 to perform monthly interlock checks and maintenance on the existing dual feed screws. A planned thermal relief valve opening is scheduled for November 9th to accomplish the above.

PREPARED BY: Steve Gibson for David Modricker, USACE

Region 3 | Mid-Atlantic Cleanup | Mid-Atlantic Superfund |EPA Home | EPA Superfund Homepage


Local Navigation


Jump to main content.